Conditional probability: Basic
At a certain college, 47% of the students are female, and 21% of the students major in finance. Furthermore, 12% of the students both are female and major in finance.

(a) What is the probability that a randomly selected finance major is female? Round your answer to 2 decimal places.
(b) What is the probability that a randomly selected female student majors in finance? Round your answer to 2 decimal places.






Answer by stanbon(75887)

Conditional probability: Basic
At a certain college, 47% of the students are female, and 21% of the students major in finance. Furthermore, 12% of the students both are female and major in finance.
(a) What is the probability that a randomly selected finance major is female? Round your answer to 2 decimal places.
---
P(female|finance major) = [P(female AND finance major)] / P[finance major]
= 0.12/0.21 = 0.5714...
=================================

(b) What is the probability that a randomly selected female student majors in finance? Round your answer to 2 decimal places.
---
P(majors in finance|female) = P(majors in fin AND female) / P(female)
= 0.12/0.47 = 0.2553..
Cheers,
Stan H.